# Partitioned Quantum Subspace Expansion

### v2

@@ 3 Results
As noted in Sec. 2.2.1, the order $O(\mathcal{R})$ of the partition $\mathcal{R}$ found by PQSE may be smaller than the maximally allowed Krylov basis order $R$. We now quantify how frequently $O(\mathcal{R})$ is smaller than $R$ and analyse the pattern of values of $O(\mathcal{R})$. Figure 4 shows the distribution of $O(\mathcal{R})$ for increasing $R$ for the spin ring example at fixed noise strength. The figure illustrates the relationship between $O(\mathcal{R})$ and $R$ as the latter increases. We see that the proportion of instances displaying early termination increases as $R$ increases and that the range of basis orders at early termination also increases. The smallest maximum allowed target Krylov basis order in this figure $R=7$, was chosen as the order at which TQSE begins to display numerical instability on average, identifiable at the minimum $\overline{\epsilon_{\text{rel}}}$ for TQSE in in Fig. 2(c). As the target basis order for PQSE increases, a large proportion of instances terminate with a basis order above this minimum. Figures 2 and 3 show that PQSE can continue increasing accuracy at basis orders above this minimum. This suggests that this implementation of PQSE is capable of detecting and avoiding numerical instabilities by choosing a stable basis sequence or terminating early.
@@
Figure 4: Distribution of the Krylov basis order of the final basis returned by PQSE $O(\mathcal{R})$ for increasing maximum allowed Krylov basis order $R$. Data for each distribution is obtained over 100 instances of a noisy 10 qubit periodic disordered Heisenberg model. Noise is drawn from Gaussian distributions width proportional to $\delta=10^{-6}$. The maximum Krylov basis order is the far right bin, coloured in red.
@@ 4 Discussion
We have presented a classical post-processing technique that, on average and for the systems under consideration, allows the QSE algorithm to be effectively utilised with a higher dimension Krylov subspace than otherwise possible. The practical implication of this is that it can enable more accurate ground state energy estimates at the cost of additional classical processing. These improvements can be substantial: enabling two orders of magnitude improvements in accuracy. Although PQSE could get closer to the true ground state energy than the other techniques, it converged slightly slower with the Krylov basis size than the other techniques considered.
The two presented examples probed the accuracy of each QSE variant in the presence of finite sampling error. In particular, we applied all three QSE approaches to estimating the ground state energy of a quasi-random spin system and a molecular Hamiltonian. We found that in both examples, PQSE obtained more accurate energy estimates on average for the majority of levels of intentionally added noise. This advantage was more pronounced for the spin system at smaller ratios of $J/h$ than for larger ratios or the molecular system. The range of minimum average errors achieved by all techniques in the spin ring example with $J/h=0.1$ and the molecular system were substantially different. This highlights that, while PQSE is able to construct sequences of eigenvalue problems with better conditioning, and thus to outperform TQSE and QSE, it is unable to overcome limitations in accuracy set by a poor choice of initial state. Improvements to the initial state, or techniques to augment the standard Krylov basis [see Eq. (7)] with improved convergence will therefore be highly valuable. The iterative construction we have proposed leaves considerable freedom to implement these improvements. For example, one could vary the criteria used to evaluate good sequences throughout the algorithm, perhaps by initially rewarding exploration before transitioning to a more targeted criterion such as the variance. Alternatively, one could consider constructing composite subspace sequences, employing the Krylov subspace when most effective.
@@
A special feature of IQAE is that the basis $\mathcal{B}_{k}$ is constructed from unitaries which do not conserve the symmetry of $H$. Therefore, if the initial state |$\ket{\phi_{0}}$ lies in a different symmetry sector to the ground state of $H$, then optimising using $\mathcal{B}_{k}$ can lower the energy of |$\ket{\phi_{0}}$ while a Krylov basis constructed with a symmetry-preserving Hamiltonian $H$ would fail. An interesting avenue for future research is to consider applying PQSE with the basis used with the IQAE algorithm, to see if the dual benefits of numerical stability and symmetry breaking can be combined. Further improvements of PQSE could include augmenting the variance criterion for choosing a good sequence of partitions, and by investigating how PQSE may be more tightly integrated with thresholding.
@@
Recent experimental demonstrations of quantum error correction [41, 42, 43, 44] provide some of the strongest evidence so far that the technological obstacles to fault tolerant quantum computation can be overcome. However, even the accuracy of early fault tolerant quantum computers may be limited by the impact of finite numerical precision and statistical noise on the output of algorithms such as QSE. Our approach suggests that this can be substantially alleviated.
@@ 6 Acknowledgements
TO acknowledges support by the EPSRC through an EPSRC iCASE studentship award in collaboration with IBM Research and the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL) under project number 20230049DR. DJ acknowledges support by the European Union’s Horizon Programme (HORIZON-CL4-2021-DIGITALEMERGING-02-10) Grant Agreement 101080085 QCFD. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work [45].
@@ Appendix A Additional numerical results
Here we provide additional data illustrating the average relative error with increasing maximum allowed Krylov basis order associated with different approaches to QSE. We present a wider range of noise strengths $\delta$ than considered in the main text. Additional results for the spin ring example are in Fig. 5 and Hydrogen molecule example in Fig. 6. A detailed description of the setup for these examples can be found in Sec. (3) in the main text. We highlight that the results shown in Fig. 3 in the main text can be recognised by taking the minimum average error across all Krylov basis orders for each noise strength $\delta>0$. Fig. 3a connects to Fig. 5 and Fig. 3b to Fig. 6.
Figure 5: Average relative error in estimate of ground state energy for QSE with a Krylov basis: unmodified (QSE), thresholded (TQSE), partitioned (PQSE) with increasing $\delta$. The target system is the 10 qubit periodic disordered Heisenberg model with $J=0.1$ and $h=1$ for increasing Gaussian noise strength.
@@
Appendix B Updating the reference state
@@ B.1 Resource cost
With the updated coefficients in hand, constructing an element of $\mathbf{S}^{\{P\}}$ using Eq. (52) requires $N_{P}$ basic steps. The same inner-product coefficients can be used to construct a single element of $\mathbf{H}^{\{P\}}$, which also requires $N_{P}$ basic steps. By making use of the Hermitian-Toeplitz structure of the $\mathbf{S}$ and $\mathbf{H}$ matrices, setting up the generalised eigenvalue problem in $\mathcal{K}_{r_{P+1}}(H,\ket{\psi})$ requires $4r_{P+1}N_{P}$ operations. Constructing all possible subspaces for a given $r_{P+1}$ and solving the generalised eigenvalue problem with cost $wr^{3}$ [33], where $w$ is a constant factor, has thus the overall cost
@@
B.2 Calculating expectation values
PQSE returns |$\ket{\psi_{P}}$ as an approximation to the ground state of $H$ in terms of expansion coefficients $\textbf{c}^{\text{opt}}_{r_{P-1}}[\ket{\phi_{P-1}}]$ relative to the Krylov basis with starting vector |$\ket{\psi_{P-1}}$. We obtain the energy of this state $E_{P}$ when solving Eq. (3) in $\mathcal{K}_{r_{P}}(H,\ket{\psi_{P-1}})$ [see Sec. 2.1 and 2.2 in the main text]. In this section we show how to compute expectation values of observables in the state |$\ket{\psi_{P}}$ in terms of expectation values in the initial state |$\ket{\phi_{0}}$. For example, this allows us to calculate the energy variance in |$\ket{\psi_{P}}$.
@@ B.2 Calculating expectation values
$\bra{\psi_{2}}\hat{O}\ket{\psi_{2}}$ $=\sum_{k_{2},l_{2}=0}^{r_{2}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt}}_{r_{2},l_{2}}\bra{\psi_{1}}H^{k_{2}}\hat{O}H^{l_{2}}\ket{\psi_{1}}\,,$ (55)
$=\sum_{k_{2},l_{2}=0}^{r_{2}-1}\sum_{k_{1},l_{1}=0}^{r_{1}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\bra{\phi_{0}}H^{k_{1}+k_{2}}\hat{O}H^{l_{1}+l_{2}}\ket{\phi_{0}}\,.$ (56)
For a given $20\leq m,m^{\prime}\leq r_{1}+r_{2}-$ 2 values of $k_{1}$, $k_{2}$ satisfying $k_{1}+k_{2}=m$ and values of $l_{1}$, $l_{2}$ satisfying $l_{1}+l_{2}=m^{\prime}$ will contribute multiple terms with the same expectation $\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}$. We can group these terms
$\bra{\psi_{2}}\hat{O}\ket{\psi_{2}}$ $=\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}\left(\mathop{\sum_{k_{2}=0}^{r_{2}-1}\sum_{k_{1},=0}^{r_{1}-1}}_{k_{1}+k_{2}=m}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}\right)\left(\mathop{\sum_{l_{2}=0}^{r_{2}-1}\sum_{l_{1}=0}^{r_{1}-1}}_{l_{1}+l_{2}=m^{\prime}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\right)\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}\,.$ (57)
$=\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}b^{*}_{m,\{r_{1},r_{2})\}}b_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}\,,$ (58)
@@
$\bra{\psi_{3}}\hat{O}\ket{\psi_{3}}$ $=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\bra{\psi_{2}}H^{k_{3}}\hat{O}H^{l_{3}}\ket{\psi_{2}}\,,$ (60)
$=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\sum_{k_{2},l_{2}=0}^{r_{2}-1}\sum_{k_{1},l_{1}=0}^{r_{1}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\bra{\phi_{0}}H^{k_{1}+k_{2}+k_{3}}\hat{O}H^{l_{1}+l_{2}+l_{3}}\ket{\phi_{0}}\,.$ (61)
$=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}b^{*}_{m,\{r_{1},r_{2}\}}b_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}H^{m+k_{3}}\hat{O}H^{m^{\prime}+l_{3}}\ket{\phi_{0}}\,.$ (62)
$=\sum_{m^{\prime\prime},m^{\prime\prime\prime}=0}^{r_{1}+r_{2}+r_{3}-3}\left(\sum_{k_{3}=0}^{r_{3}-1}\sum_{m=0}^{r_{1}+r_{2}-2}c^{\text{opt},*}_{k_{3}}b^{*}_{m,\{r_{1},r_{2}\}}\right)\left(\sum_{l_{3}=0}^{r_{3}-1}\sum_{m^{\prime}=0}^{r_{1}+r_{2}-2}c^{\text{opt}}_{l_{3}}b_{m^{\prime},\{r_{1},r_{2}\}}\right)\bra{\phi_{0}}H^{m^{\prime\prime}}\hat{O}H^{m^{\prime\prime\prime}}\ket{\phi_{0}}\,.$ (63)
$=\sum_{m^{\prime\prime},m^{\prime\prime\prime}=0}^{r_{1}+r_{2}+r_{3}-3}b^{*}_{m^{\prime\prime},\{r_{1},r_{2},r_{3}\}}b_{m^{\prime\prime\prime},\{r_{1},r_{2},r_{3}\}}\bra{\phi_{0}}H^{m^{\prime\prime}}\hat{O}H^{m^{\prime\prime\prime}}\ket{\phi_{0}}\,,$ (64)

### v3

@@ 3 Results
We now present a brief investigation into the typical behaviour of PQSE. As noted in Sec. 2.2.1, the order $O(\mathcal{R})$ of the partition $\mathcal{R}$ found by PQSE may be smaller than the maximally allowed Krylov basis order $R$. We now quantify how frequently $O(\mathcal{R})$ is smaller than $R$ and analyse the pattern of values of $O(\mathcal{R})$. Figure 4 shows the distribution of $O(\mathcal{R})$ for increasing $R$ for the spin ring example at fixed noise strength. The figure illustrates the relationship between $O(\mathcal{R})$ and $R$ as the latter increases. We see that the proportion of instances displaying early termination increases as $R$ increases and that the range of basis orders at early termination also increases. As the target basis order for PQSE increases, a large proportion of instances terminate with a basis order above this minimum. Figures 2 and 3 show that PQSE can continue increasing accuracy at basis orders above this minimum. This suggests that this implementation of PQSE is capable of detecting and avoiding numerical instabilities by choosing a stable basis sequence or terminating early. An extension to these results can be found in Appendix D, where we provide overview of the typical basis-size sequences chosen by PQSE for the spin ring example with $\delta=10^{-6}$.
Additional results for $\overline{\epsilon_{\text{rel}}}$ as a function of $R$ for a range of noise strengths are provided in Appendix A and show similar qualitative behaviour to the $\delta=10^{-6}$ case. Furthermore, in Appendix A we also recompute Figs. 3(a) and (b) for the Heisenberg model and $\text{H}_{6}$ examples using a PQSE criterion which avoids using the matrix element $\bra{\phi_{0}}H^{2R}\ket{\phi_{0}}$, finding performance to be very similar in both cases. Finally, in order to present results in a more near term relevant setting, as well as providing evidence that PQSE generalize to other Krylov basis generation methods, we present the results of our numerical experiments with a Krylov basis generated through real time evolution in Appendix (B). We see comparative performance between PQSE and parameter-optimized TQSE, motivating application of PQSE in near-term experiments.
@@
Figure 4: Distribution of the Krylov basis order of the final basis returned by PQSE $O(\mathcal{R})$ for increasing maximum allowed Krylov basis order $R$. Data for each distribution is obtained over 200 instances of a noisy 10 qubit periodic disordered Heisenberg model. Noise is drawn from Gaussian distributions width proportional to $\delta=10^{-6}$. The maximum Krylov basis order is the far right bin, coloured in red.
@@ 4 Discussion
We have presented a classical post-processing technique that allows the QSE algorithm to be effectively utilised with a higher dimension Krylov subspace without requiring problem and noise dependent parameter tuning. The practical implication of this is that it can enable more accurate ground state energy estimates. In addition to competitive performance with the state of the art, PQSE can offer order of magnitude improvements in accuracy.
The two presented examples probed the accuracy of each QSE variant in the presence of finite sampling error. In particular, we applied all three QSE approaches to estimating the ground state energy of a quasi-random spin system and a molecular Hamiltonian. We found that for the spin ring example at smaller ratios of $J/h$ and the molecular example PQSE either performed better or comparatively to TQSE with a optimized parameterisation. As the TQSE threshold was chosen by optimising with respect to the true ground state energy, we argue that this is close to the best-case performance for TQSE. That PQSE remains competitive under these conditions is particularly noteworthy. For larger ratios of $J/h$, TQSE was more accurate than PQSE. Obtaining a clearer understanding of this discrepancy, perhaps whether it is due to sensitivity to the choice of |$\ket{\phi_{0}}$, is an interesting starting point for future work. The range of minimum average errors achieved by all techniques in the spin ring example with $J/h=0.1$ and the molecular system were substantially different. This highlights that, while PQSE is able to construct sequences of eigenvalue problems with better conditioning, it is unable to overcome limitations in accuracy set by a poor choice of initial state. Improvements to the initial state, or techniques to augment the standard Krylov basis [see Eq. (7)] with improved convergence will therefore be highly valuable. Additionally, for both PQSE and TQSE, we find that the variational principle constraining the trial solution energy to be less than the true ground state energy of our target Hamiltonian is broken. Ensuring this principle holds would be a useful improvement for future work. The iterative construction we have proposed leaves considerable freedom to implement these improvements. For example, one could vary the criteria used to evaluate good sequences throughout the algorithm, perhaps by initially rewarding exploration before transitioning to a more targeted criterion such as the variance. Alternatively, one could consider constructing composite subspace sequences, employing the Krylov subspace when most effective. These improvements could draw from classical iterative Krylov approaches such as the Restarted Arnoldi algorithm [44], which also uses reference state re-initialisation, and the Davidson algorithm [45], which has already seen development in a quantum context [46, 47] and involves carefully growing a Krylov basis through a process of computing residual vectors and applying a pre-conditioner.
@@
A special feature of IQAE is that the basis $\mathcal{B}_{k}$ is constructed from unitaries which do not conserve the symmetry of $H$. Therefore, if the initial state |$\ket{\phi_{0}}$ lies in a different symmetry sector to the ground state of $H$, then optimising using $\mathcal{B}_{k}$ can lower the energy of |$\ket{\phi_{0}}$ while a Krylov basis constructed with a symmetry-preserving Hamiltonian $H$ would fail. An interesting avenue for future research is to consider applying PQSE with the basis used with the IQAE algorithm, to see if the dual benefits of numerical stability and symmetry breaking can be combined. A similar direction could be using classical shadow data, which can be applied to enhance the performance of subspace methods with a poor initial state [48]. Further improvements of PQSE could include augmenting the variance criterion for choosing a good sequence of partitions, and by investigating how PQSE may be more tightly integrated with thresholding.
@@
Recent experimental demonstrations of quantum error correction [49, 50, 51, 52] provide some of the strongest evidence so far that the technological obstacles to fault tolerant quantum computation can be overcome. However, even the accuracy of early fault tolerant quantum computers may be limited by the impact of finite numerical precision and statistical noise on the output of algorithms such as QSE. Our approach suggests that this can be strongly alleviated in a parameter-free way.
@@ 6 Acknowledgements
T. O. acknowledges support by the EPSRC through an EPSRC iCASE studentship award in collaboration with IBM Research (EP/W522211/1) and the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL) under project number 20230049DR. D. J. acknowledges support by the European Union’s Horizon Programme (HORIZON-CL4-2021-DIGITALEMERGING-02-10) Grant Agreement 101080085 QCFD. D. J. acknowledges support from the Hamburg Quantum Computing Initiative (HQIC) project EFRE. The project is co-financed by ERDF of the European Union and by “Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB) The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work [53].
@@ Appendix A Additional numerical results
Here we provide additional data illustrating the average relative error with increasing maximum allowed Krylov basis order associated with different approaches to QSE. We present a wider range of noise strengths $\delta$ than considered in the main text. Additional results for the spin ring example are in Fig. 5 and Hydrogen chain example in Fig. 6. A detailed description of the setup for these examples can be found in Sec. (3) in the main text. We highlight that the results shown in Fig. 3 in the main text can be recognised by taking the minimum average error across all Krylov basis orders for each noise strength $\delta>0$. Fig. 3a connects to Fig. 5 and Fig. 3b to Fig. 6.
Additionally, we present the minimum average error $\xi$ with increasing noise strength when PQSE is using an alternative basis selection criterion. Here, once PQSE has constructed a state involving a terminating Hamiltonian power: $H^{R-1}$, it will use the state’s energy squared, computed using the procedure in Appendix (C.3), instead of the energy variance as a metric to compare with other possible solutions. We present the corresponding results in Figure 7, finding that the performance of PQSE here closely resembles that in the main text Figure (3).
Figure 5: Average relative error in estimate of ground state energy for QSE with a Krylov basis generated using Hamiltonian powers: unmodified (QSE), thresholded (TQSE), partitioned (PQSE) with increasing $\delta$. The target system is the 10 qubit periodic disordered Heisenberg model with $J=0.1$ and $h=1$ for increasing Gaussian noise strength.
Figure 6: Average relative error in estimate of ground state energy for QSE with a Krylov basis generated using Hamiltonian powers: unmodified (QSE), thresholded (TQSE), partitioned (PQSE) with increasing $\delta$. The target system is H6 in a STO-3G basis with a parity mapping applied in equilibrium configuration.
Figure 7: Minimum average relative error $\xi$ in the estimate of ground state energy for QSE with a Krylov basis generated using Hamiltonian powers: unmodified (QSE), thresholded (TQSE), partitioned (PQSE). Here PQSE uses a more economical basis selection criterion, comparing solution energy instead of energy variance at a terminating Krylov basis order. (a) $\xi$ as a function of noise strength $\delta$ for a 10 qubit periodic disordered Heisenberg model with $J/h=0.1$. (b) $\xi$ as a function of noise strength $\delta$ for H6 in a STO-3G basis with a parity mapping applied in equilibrium configuration. Error bars show the standard error.
Appendix B Generating a Krylov Basis through Real Time Evolution
In this section we present the results for our spin ring and molecular examples using a Krylov basis generated in by real time evolution. This approach is more suited to near term quantum computers as it typically involves shallower quantum circuits than those needed to generate a Krylov basis with Hamiltonian powers. The subspace is generated by real time evolution (RTE)
$\tilde{\mathcal{K}}_{R}(H,\ket{\phi_{0}})=\text{span}\{e^{-\mathrm{i}Ht_{j}}\ket{\phi_{0}}\}_{j=0}^{R-1},$ (23)
@@ Appendix B Generating a Krylov Basis through Real Time Evolution
$H_{ij}$ $=\braket{\psi_{i}}{H}{\psi_{j}}\,,$ (24a)
$S_{ij}$ $=\braket{\psi_{i}}{\psi_{j}}\,,$ (24b)
$i,j$$\in\{0,1,\ldots,R-1\},$ (24c)
where |$\ket{\psi_{j}}=e^{-\mathrm{i}j\text{d}tH}\ket{\phi_{0}}$.
We consider the same numerical setup as described in the main text. For TQSE we follow the same parameter scan procedure. We highlight two differences to the main text numerical experiments necessitated by the change of basis
  1. 1.
    Finite shot noise implementation. With the RTE basis we have to estimate the finite shot error involved in computing matrix elements of the form $\bra{\psi_{i}}H\ket{\psi_{j}}$ and $\braket{\psi_{i}\lvert\psi_{j}}$. We add noise in the same way as the power basis experiments by assuming the measurement error in estimating $\bra{\psi_{i}}H\ket{\psi_{j}}$ is well approximated by the error in estimating $\bra{\psi_{j}}H\ket{\psi_{j}}$. This is motivated by existing procedures for estimating off-diagonal matrix elements e.g. see Section 2.1 of [54]. For $S_{ij}=\braket{\psi_{i}\lvert\psi_{j}}$ we draw noise from a distribution of width $\delta$ instead of $\delta\sqrt{\text{Var}[S_{ij}]}$.
  2. 2.
    Variance criterion matrix elements. For consistency with the rest of the paper we use the variance minimization criterion for PQSE again. However, we note that we no longer get the matrix elements required to compute solution energy variance “for free”, as $\bra{\psi_{i}}H^{2}\ket{\psi_{j}}$ is not used in QSE. We add noise in the same way to the other matrix elements.
We now present the results using a RTE basis. In Fig. 8 we include the minimum average error at each noise strength for both numerical examples and each QSE approach. In Fig. 8(a), for the spin ring example, we see that PQSE and TQSE have very similar accuracy. TQSE has a slight advantage at $\delta=10^{-10}$, which increases up to around half an order of magnitude improvement at $\delta=10^{-1}$. In Fig. 8(b), for the H6 example, we see that PQSE and TQSE have similar accuracy. TQSE has an advantage of approximately half an order of magnitude for all noise strengths. The full set of results used to produce these plots are included in Figs. 9 and 10 for the spin ring and molecular examples respectively. These results highlight that PQSE is also competitive with parameter-optimized TQSE for the RTE-generated Krylov basis. This is notable since the TQSE parameter optimization used knowledge of the exact ground state energy. Although these results worked with exact time evolution and in practice some approximation error will be present, we consider this as positive motivation for applying PQSE to the output of near-term experiments.
(a)Refer to caption
(b)Refer to caption
Figure 8: Minimum average relative error $\xi$ in the estimate of ground state energy for QSE with a Krylov basis generated using real time evolution: unmodified (QSE), thresholded (TQSE), partitioned (PQSE). (a) $\xi$ as a function of noise strength $\delta$ for a 10 qubit periodic disordered Heisenberg model with $J/h=0.1$. (b) $\xi$ as a function of noise strength $\delta$ for H6 in a STO-3G basis with a parity mapping applied in equilibrium configuration. Error bars show the standard error.
Figure 9: Average relative error in estimate of ground state energy for QSE with a Krylov basis generated using real time evolution: unmodified (QSE), thresholded (TQSE), partitioned (PQSE) with increasing $\delta$. The target system is the 10 qubit periodic disordered Heisenberg model with $J=0.1$ and $h=1$ for increasing Gaussian noise strength.
Figure 10: Average relative error in estimate of ground state energy for QSE with a Krylov basis using real time evolution: unmodified (QSE), thresholded (TQSE), partitioned (PQSE) with increasing $\delta$. The target system is H6 in a STO-3G basis with a parity mapping applied in equilibrium configuration.
Appendix C Updating the reference state
C.1 Hamiltonian power Krylov basis
@@ C.2 Resource cost
With the updated coefficients in hand, constructing an element of $\mathbf{S}^{\{P\}}$ using Eq. (45) requires $N_{P}$ basic steps. The same inner-product coefficients can be used to construct a single element of $\mathbf{H}^{\{P\}}$, which also requires $N_{P}$ basic steps. By making use of the Hankel structure of the $\mathbf{S}$ and $\mathbf{H}$ matrices, setting up the generalised eigenvalue problem in $\mathcal{K}_{r_{P+1}}(H,\ket{\psi})$ requires $4r_{P+1}N_{P}$ operations. Constructing all possible subspaces for a given $r_{P+1}$ and solving the generalised eigenvalue problem with cost $wr^{3}$ [36], where $w$ is a constant factor, has thus the overall cost
@@
C.3 Calculating expectation values for a general Krylov basis
PQSE returns |$\ket{\psi_{P}}$ as an approximation to the ground state of $H$ in terms of expansion coefficients $\textbf{c}^{\text{opt}}_{r_{P-1}}[\ket{\phi_{P-1}}]$ relative to the Krylov basis with starting vector |$\ket{\psi_{P-1}}$. We obtain the energy of this state $E_{P}$ when solving Eq. (3) in $\mathcal{K}_{r_{P}}(A,\ket{\psi_{P-1}})$ [see Sec. 2.1 and 2.2 in the main text]. In this section we show how to compute expectation values of observables in the state |$\ket{\psi_{P}}$ in terms of expectation values in the initial state |$\ket{\phi_{0}}$ and a Krylov basis generating operator $A$. For example, this allows us to calculate the energy variance in |$\ket{\psi_{P}}$.
@@ C.3 Calculating expectation values for a general Krylov basis
$\bra{\psi_{2}}\hat{O}\ket{\psi_{2}}$ $=\sum_{k_{2},l_{2}=0}^{r_{2}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt}}_{r_{2},l_{2}}\bra{\psi_{1}}(A^{k_{2}})^{\dagger}\hat{O}A^{l_{2}}\ket{\psi_{1}}\,,$ (48)
$=\sum_{k_{2},l_{2}=0}^{r_{2}-1}\sum_{k_{1},l_{1}=0}^{r_{1}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\bra{\phi_{0}}(A^{k_{1}+k_{2}})^{\dagger}\hat{O}A^{l_{1}+l_{2}}\ket{\phi_{0}}\,.$ (49)
For a given $20\leq m,m^{\prime}\leq r_{1}+r_{2}-$ 2 values of $k_{1}$, $k_{2}$ satisfying $k_{1}+k_{2}=m$ and values of $l_{1}$, $l_{2}$ satisfying $l_{1}+l_{2}=m^{\prime}$ will contribute multiple terms with the same expectation $\bra{\phi_{0}}(A^{m})^{\dagger}\hat{O}A^{m^{\prime}}\ket{\phi_{0}}$. We can group these terms
$\bra{\psi_{2}}\hat{O}\ket{\psi_{2}}$ $=\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}\left(\mathop{\sum_{k_{2}=0}^{r_{2}-1}\sum_{k_{1},=0}^{r_{1}-1}}_{k_{1}+k_{2}=m}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}\right)\left(\mathop{\sum_{l_{2}=0}^{r_{2}-1}\sum_{l_{1}=0}^{r_{1}-1}}_{l_{1}+l_{2}=m^{\prime}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\right)\bra{\phi_{0}}(A^{m})^{\dagger}\hat{O}A^{m^{\prime}}\ket{\phi_{0}}\,.$ (50)
$=\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}b^{*}_{m,\{r_{1},r_{2})\}}b_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}(A^{m})^{\dagger}\hat{O}A^{m^{\prime}}\ket{\phi_{0}}\,,$ (51)
@@
$\bra{\psi_{3}}\hat{O}\ket{\psi_{3}}$ $=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\bra{\psi_{2}}(A^{k_{3}})^{\dagger}\hat{O}A^{l_{3}}\ket{\psi_{2}}\,,$ (53)
$=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\sum_{k_{2},l_{2}=0}^{r_{2}-1}\sum_{k_{1},l_{1}=0}^{r_{1}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\bra{\phi_{0}}(A^{k_{1}+k_{2}+k_{3}})^{\dagger}\hat{O}A^{l_{1}+l_{2}+l_{3}}\ket{\phi_{0}}\,.$ (54)
$=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}b^{*}_{m,\{r_{1},r_{2}\}}b_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}(A^{m+k_{3}})^{\dagger}\hat{O}A^{m^{\prime}+l_{3}}\ket{\phi_{0}}\,.$ (55)
$=\sum_{m^{\prime\prime},m^{\prime\prime\prime}=0}^{r_{1}+r_{2}+r_{3}-3}\left(\sum_{k_{3}=0}^{r_{3}-1}\sum_{m=0}^{r_{1}+r_{2}-2}c^{\text{opt},*}_{k_{3}}b^{*}_{m,\{r_{1},r_{2}\}}\right)\left(\sum_{l_{3}=0}^{r_{3}-1}\sum_{m^{\prime}=0}^{r_{1}+r_{2}-2}c^{\text{opt}}_{l_{3}}b_{m^{\prime},\{r_{1},r_{2}\}}\right)\bra{\phi_{0}}(A^{m^{\prime\prime}})^{\dagger}\hat{O}A^{m^{\prime\prime\prime}}\ket{\phi_{0}}\,.$ (56)
$=\sum_{m^{\prime\prime},m^{\prime\prime\prime}=0}^{r_{1}+r_{2}+r_{3}-3}b^{*}_{m^{\prime\prime},\{r_{1},r_{2},r_{3}\}}b_{m^{\prime\prime\prime},\{r_{1},r_{2},r_{3}\}}\bra{\phi_{0}}(A^{m^{\prime\prime}})^{\dagger}\hat{O}A^{m^{\prime\prime\prime}}\ket{\phi_{0}}\,,$ (57)